Project

General

Profile

NDOS Commissioning Meeting 2012/03/12

Attendance: Jon, Peter, Tomas, Martin, Leon, Mark, Mat, Andrew

Agenda:

  1. Beam and Detector Update
  2. DataCheck Report
  3. Installation Plans for the week
  4. DAQ Update and Plans for the week
  5. DCS Update and Plans for the week
  6. AOB

Beam and Detector Report

  • Last week's running generally went ok
  • Last Friday, TDU in MINOS exhibited strange behavior (could not lock onto GPS signal). This was replaced by TDU from teststand. The failed TDU is still in MINOS for study.
  • Today 4 new DCMs were installed along with a new TDU. The new DCMs and TDU have new hardware and firmware and are incompatible with the old DCMs and TDU. A run was successfully started earlier today with the old hardware only; experts are working now on getting the system up and running with the new hardware.

DataCheck Report (Jan):

  • No report

Installation plans for this week

  • Leon: plan to install APDs later this week. All parts should be here by Wed. afternoon.
  • Mat: plan is to extend fiducial in both views upstream of currently instrumented

DAQ Update and Plans for next week

  • Peter: Kurt did a DDS test of multicast scheme, would be good to continue these studies after the dust settles this week.

DCS Update and Plans for next week

  • No report.
  • Mat: we will need 38 channels of analog read-back for dry gas system; where should we run the labview? Need to get another monitor set up on the windows machine. Chad is helping Mat write the labview code.

AOB

  • Mat: Carl R. took 2 cable harnesses off of the detector and replace the one cable from each that was attached to a DCM by a single (pre-tested) cable.